OpenStudy (anonymous):

Find an exact value sin(19pi/12)

OpenStudy (anonymous):

\(\frac{19\pi}{12}\) is half of \(\frac{19\pi}{6}\) use the "half angle" formula

OpenStudy (anonymous):

is that a 10x/12

OpenStudy (anonymous):

I still don't see what I need to do

OpenStudy (anonymous):

\[\bf \sin(x)= \pm \sqrt{\frac{ 1-\cos(2x) }{ 2 }}\]We use this Half-Angle formula to evaluate what you're looking for. We can make the x in sin(x) x/2 and make the 2x in cos(2x) just x by dividing each by 2.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sin(19pi/12)=+-sqrt1-cos(19pi/6)/2

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

what do I do after

OpenStudy (anonymous):

I have to finish this tomorrow

OpenStudy (anonymous):

Now evaluate the expression under the square root. @torobi

OpenStudy (anonymous):

Ok I figured it out so I'm going to put it up here for the next person that searches this question \[\sin(\frac{ 19\pi }{ 12})\] \[\sin(\frac{ 9\pi }{ 12 }+\frac{ 10\pi }{ 12 })\] because these 2 numbers add up to 19pi and when reduced can be found on the unit cirlce \[\sin (\frac{ 3\pi }{ 4 }+\frac{ 5\pi }{ 6 })\] Use the sin sum formula and the rest should be easy